OpenStudy (anonymous):

PLEAS HELP IM BEGGING! Solve 3x2 = -12x - 15.

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

is the first part of the equation (3x)(2)?

OpenStudy (cwrw238):

3x^2 +12x + 15 = 0 divide through by 3:- 3(x^2 + 4x + 5) = 0

OpenStudy (anonymous):

okay then what

OpenStudy (cwrw238):

you can use the quadratic formula to solve this - it will not factor

OpenStudy (anonymous):

i got -4 +/- i

OpenStudy (cwrw238):

x = [-4 +/- sqrt(4^4 -4*5)] / 2 = -4 +/- sqrt(-4) / 2 = [-4 +/- 2i] / 2 = -2 +/- i

OpenStudy (anonymous):

oh wow okay i get it thank you so much!

OpenStudy (cwrw238):

yw